# Simulating Polynomial-Time Nondeterministic Turing Machines via Nondeterministic Turing Machines

### v19

@@ Abstract.
We prove in this paper that there is a language $L_{s}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$ machines. Then we further show that $L_{s}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The main techniques used in this paper are simulation and the novel new techniques developed in the author’s recent work. Further, if there exists some oracle $A$ such that $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mystery lies behind it and show that if $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all ${\rm co}\mathcal{NP}^{A}$ machines is not enumerable, thus showing that the simulation techniques are not applicable for the first half of the whole step to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$. Our main result has profound implications, such as $\mathcal{P}\neq\mathcal{NP}$ and a lower bounds result for Frege proof systems (i.e., no Frege proof systems can be polynomially bounded), etc.
@@ 1. Introduction
If $\mathcal{NP}={\rm co}\mathcal{NP}$, then for every problem where “yes” answers are efficiently verifiable, “no” answers would also be efficiently verifiable. This would be surprising because many problems, such as Satisfiability [Coo71, Lev73] or Graph non-isomorphism [GJ79], seem to lack obvious short certificates for “no” instances. It is also widely believed that $\mathcal{NP}={\rm co}\mathcal{NP}$ would imply significant consequences for the Polynomial hierarchy ($PH$) [Sto77, A6], a layered structure of complexity classes built on $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$. Specifically, if $\mathcal{NP}={\rm co}\mathcal{NP}$, then $PH$ would collapse to $\mathcal{NP}$, i.e., $PH=\Sigma_{1}^{P}=\prod_{1}^{P}=\mathcal{NP}$; see e.g. Theorem 5.4 in [AB09]. This contradicts the widely believed assumption that $PH$ is infinitely extended. Note that this conjecture is not only one of the holy grail problems of theoretical computer science but also the core key to understanding the essence of computing.
@@
Our motivation in this paper is that, in the author’s recent work [Lin21], we left this outstanding and long-standing open conjecture $\mathcal{NP}\overset{?}{=}{\rm co}\mathcal{NP}$ untouched. In the present paper, we try to resolve this important open conjecture.
@@ 1.1. Our Contributions
Corollary 1.3.
$\mathcal{P}\neq\mathcal{NP}$.
Moreover, let $\mathcal{NEXP}$ and ${\rm co}\mathcal{NEXP}$ be the complexity classes defined by
@@ 1.2. Background and Prior Work
where $L$ is a decision problem, and $\overline{L}$ is the complement of $L$ (i.e., $\overline{L}=\Sigma^{*}\setminus L$ with the assumption that the language $L$ is over the alphabet $\Sigma$). That is to say, ${\rm co}\mathcal{NP}$ is the complement of $\mathcal{NP}$. Note that the complement of a decision problem $L$ is defined as the decision problem whose answer is “yes” whenever the input is a “no” input of $L$, and vice versa. In other words, for instance, according to Wikipedia [A3], ${\rm co}\mathcal{NP}$ is the set of decision problems where there exists a polynomial $p(n)$ and a polynomial-time bounded Turing machine $M$ such that for every instance $x$, $x$ is a no-instance if and only if for some possible certificate $c$ of length bounded by $p(n)$, the Turing machine $M$ accepts the pair $(x,c)$. To the best of our knowledge, the complexity class ${\rm co}\mathcal{NP}$ was introduced for the first time by Meyer and Stockmeyer [MS72] with the name “$\prod_{1}^{P}$,” and Stockmeyer wrote a full paper on the polynomial hierarchy, which also uses the notation ${\rm co}\mathcal{NP}$; see e.g. [Sto77].
The $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem is deeply connected to the field of proof complexity. Here are some historical notes: In 1979, Cook and Reckhow [CR79] introduced a general definition of a propositional proof system and related it to mainstream complexity theory by pointing out that such a system exists in which all tautologies have polynomial-length proofs if and only if the two complexity classes $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ coincide; see e.g. [CN10]. We refer the reader to the reference [Coo00] for the importance of this research field and to the reference [Kra95] for the motivation of the development of this rich theory. Apart from those, Chapter $10$ of [Pap94] also contains the introductions of the importance of the problem
@@
Likewise, propositional proof complexity is an area of study that has seen a rapid development over the last two decades, which plays as important a role in the theory of feasible proofs as the role played by the complexity of Boolean circuits in the theory of efficient computations; see e.g. the celebrated work [Raz15] by Razborov. In most cases, according to [Raz15], the basic question of propositional proof complexity can be reduced to that given a mathematical statement encoded as a propositional tautology $\psi$ and a class of admissible mathematical proofs formalized as a propositional proof system $P$, what is the minimal possible complexity of a $P$-proof of $\psi$? In other words, propositional proof complexity aims to understand and analyze the computational resources required to prove propositional tautologies, in the same way that circuit complexity studies the resources required to compute Boolean functions. In light of the point of view given in [Raz15], the task of proving lower bounds for strong proof systems like Frege or Extended Frege modulo any hardness assumption in the purely computational world may be almost as interesting; see e.g., [Raz15]. It is worth noting that it has been observed by Razborov [Raz15] that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ implies lower bounds for any propositional proof system. In addition, see e.g., [Raz03] for interesting directions in propositional proof complexity.
@@
Let us return back to the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem for this moment. As a matter of fact, the so-called $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem is the central problem in proof complexity; see e.g. the excellent survey [Kra19]. It formalizes the question of whether or not there are efficient ways to prove the negative cases in our and in many other similar examples. Moreover, the ultimate goal of proof complexity is to show that there is no universal propositional proof system allowing for efficient proofs of all tautologies, which is equivalent to showing that the computational complexity class $\mathcal{NP}$ is not closed under the complement; see e.g. [Kra19].
Particularly, in 1974, Cook and Reckhow showed in [CR74] that there exists a super proof system if and only if $\mathcal{NP}$ is closed under complement; that is, if and only if $\mathcal{NP}={\rm co}\mathcal{NP}$. It is, of course, open whether any superproof system exists. The above result has led to what is sometimes called “Cook’s program” for proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$: prove superpolynomial lower bounds for proof lengths in stronger and stronger proposition proof systems until they are established for all abstract proof systems. Cook’s program, which also can be seen as a program to prove $\mathcal{P}\neq\mathcal{NP}$, is an attractive and plausible approach; unfortunately, it has turned out to be quite hard to establish superpolynomial lower bounds on common proof systems; see, e.g., the excellent survey on proof complexity [Bus12].
@@ 2. Preliminaries
Let $\mathbb{N}_{1}=\{1,2,3,\cdots\}$, where $+\infty\notin\mathbb{N}_{1}$, i.e., the set of all positive integers. We also denote by $\mathbb{N}$ the set of natural numbers, i.e.,
@@
A ${\rm co}\mathcal{NP}$ machine $M$ is exactly a polynomial-time nondeterministic Turing machine $M$ working under the “there exists no accepting path” condition. We define the language accepted by the ${\rm co}\mathcal{NP}$ machine $M$ to be $\overline{L}(M)$. 222 In short, there is an interesting relationship: ${\rm co}\mathcal{NP}$ machine $M$ accepts the language $\overline{L}(M)$, but polynomial-time nondeterministic Turing machine $M$ rejects $\overline{L}(M)$; polynomial-time nondeterministic Turing machine $M$ accepts the language $L(M)$, but ${\rm co}\mathcal{NP}$ machine $M$ rejects $L(M)$. Namely, for any input $w$, $w\in\overline{L}(M)$ if and only if polynomial-time nondeterministic Turing machine $M$ rejects $w$.
@@ 3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
4. Simulating of All Polynomial-time Nondeterministic Turing Machines
In computational complexity, the most basic approach to show hierarchy theorems uses simulation and diagonalization, because the simulation and diagonalization technique [Tur37, HS65] is a standard method to prove lower bounds on uniform computing models (i.e., the Turing machine model); see for example [AHU74]. These techniques work well in deterministic time and space measures; see e.g., [For00, FS07]. However, they do not work for nondeterministic time, which is not known to be closed under complement at present, i.e., one of the main issues discussed in this paper; hence it is unclear how to define a nondeterministic machine that “does the opposite”; see e.g., [FS07].
For nondeterministic time, we can still do a simulation but can no longer negate the answer directly. In this case, we apply the lazy diagonalization to show the nondeterministic time hierarchy theorem; see e.g. [AB09]. It is worth noting that Fortnow [For11] developed a much more elegant and simple style of diagonalization to show nondeterministic time hierarchy; see also [FS10]. Generally, lazy diagonalization [Zak83, AB09, SFM78] is a clever application of the standard diagonalization technique [Tur37, HS65, AHU74]. The basic strategy of lazy diagonalization is that for any $n\leq j<m=n^{t(n)}$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $1^{j+1}$, accepting if $M_{i}$ accepts and rejecting if $M_{i}$ rejects. When $j=m$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $x$ deterministically, accepting if $M_{i}$ rejects and rejecting if $M_{i}$ accepts. Since $m=n^{t(n)}$, $M$ has enough time to perform the trivial exponential-time deterministic simulation of $M_{i}$ on input $1^{n}$; what we are doing is deferring the diagonalization step by linking $M$ and $M_{i}$ together on larger and larger inputs until $M$ has an input large enough that it can actually “do the opposite” deterministically; see e.g. [FS07] for details.
@@ 5. Proving That $L_{s}\in\mathcal{NP}$
For each fixed $k\in\mathbb{N}_{1}$, it holds that
@@ 6. Breaking the “Relativization Barrier”
Let us suppose that $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$. In this case, to break the Relativization Barrier or to explore the mystery behind the implications of Baker, Gill, and Solovay’s result [BGS75], let us further make the following rational assumptions:
@@ A Frege proof system $F$ is an inference system for propositional logic based on
The axiom schemes of a Frege proof system is the following (see e.g. [Bus99]):
@@ 8.1. Lower Bounds for Frege Systems
Initially, propositional proof complexity has been primarily concerned with proving lower bounds (even conditional) for the length of proofs in propositional proof systems, which is extremely interesting and well justified in its own right, with the ultimate goal of settling whether $\mathcal{NP}={\rm co}\mathcal{NP}$ (see e.g. [CR79]). In fact, as mentioned in Section 1, such research directions are called Cook’s Program for separating $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ (see e.g. [Bus12]): Prove superpolynomial lower bounds for proof lengths in stronger and stronger propositional proof systems until they are established for all abstract proof systems. However, our approach is “to do the opposite,” that is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems.
@@ 9. Open Problems
In addition, does the complexity class $\mathcal{BQP}$ (see page 213 of [AB09]) contain the problem Satisfiability? Note that by footnote $1$, we have $\mathcal{BQP}\neq\mathcal{NP}$, but this can not exclude that $\mathcal{NP}\subsetneqq\mathcal{BQP}$. If one can show that the problem of Satisfiability is not in $\mathcal{BQP}$, then the relationship between $\mathcal{BQP}$ and $\mathcal{NP}$ will were clear. Currently, we have no additional information about this problem. We tried to think about it but also failed. We remark that this is also a fascinating open problem in complexity theory.

### v20

@@ Abstract.
We prove in this paper that there is a language $L_{s}$ accepted by some nondeterministic Turing machine that runs within time $O(n^{k})$ for any positive integer $k\in\mathbb{N}_{1}$ but not by any ${\rm co}\mathcal{NP}$ machines. Then we further show that $L_{s}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The main techniques used in this paper are simulation and the novel new techniques developed in the author’s recent work. Our main result has profound implications, such as $\mathcal{P}\neq\mathcal{NP}$, etc. Further, if there exists some oracle $A$ such that $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mystery lies behind it and show that if $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all ${\rm co}\mathcal{NP}^{A}$ machines is not enumerable, thus showing that the simulation techniques are not applicable for the first half of the whole step to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$. Finally, a lower bounds result for Frege proof systems is presented (i.e., no Frege proof systems can be polynomially bounded).
@@ 1. Introduction
If $\mathcal{NP}={\rm co}\mathcal{NP}$, then for every problem where “yes” answers are efficiently verifiable, “no” answers would also be efficiently verifiable. This would be surprising because many problems, such as Satisfiability [Coo71, Lev73] or Graph non-isomorphism [GJ79], seem to lack obvious short certificates for “no” instances. It is also widely believed that $\mathcal{NP}={\rm co}\mathcal{NP}$ would imply significant consequences for the Polynomial hierarchy ($PH$) [Sto77, A6], a layered structure of complexity classes built on $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$. Specifically, if $\mathcal{NP}={\rm co}\mathcal{NP}$, then $PH$ would collapse to $\mathcal{NP}$, i.e., $PH=\Sigma_{1}^{P}=\prod_{1}^{P}=\mathcal{NP}$; see e.g., Theorem 5.4 in [AB09]. This contradicts the widely believed assumption that $PH$ is infinitely extended. Note that this conjecture is not only one of the holy grail problems of theoretical computer science but also the core key to understanding the essence of computing.
@@
Our motivation in this paper is that, in the author’s recent work [Lin21], where we presented a proof that $\mathcal{P}\neq\mathcal{NP}$, we were completely at a loss and didn’t know how to prove $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, so we left this outstanding and long-standing conjecture of $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ untouched. In the present paper, we try to resolve this important open conjecture.
@@ 1.1. Our Contributions
Corollary 1.2.
$\mathcal{P}\neq\mathcal{NP}$ [^2] So, from this paper, we know that resolving the $\mathcal{P}\overset{?}{=}\mathcal{NP}$ problem does not necessarily require diagonalization techniques.
@@
Furthermore, let $\mathcal{NEXP}$ and ${\rm co}\mathcal{NEXP}$ be the complexity classes defined by
@@ 1.2. Background and Prior Work
where $L$ is a decision problem, and $\overline{L}$ is the complement of $L$ (i.e., $\overline{L}=\Sigma^{*}\setminus L$ with the assumption that the language $L$ is over the alphabet $\Sigma$). That is to say, ${\rm co}\mathcal{NP}$ is the complement of $\mathcal{NP}$. Note that the complement of a decision problem $L$ is defined as the decision problem whose answer is “yes” whenever the input is a “no” input of $L$, and vice versa. In other words, for instance, according to Wikipedia [A3], ${\rm co}\mathcal{NP}$ is the set of decision problems where there exists a polynomial $p(n)$ and a polynomial-time bounded Turing machine $M$ such that for every instance $x$, $x$ is a no-instance if and only if for some possible certificate $c$ of length bounded by $p(n)$, the Turing machine $M$ accepts the pair $(x,c)$. To the best of our knowledge, the complexity class ${\rm co}\mathcal{NP}$ was introduced for the first time by Meyer and Stockmeyer [MS72] with the name “$\prod_{1}^{P}$,” and Stockmeyer wrote a full paper on the polynomial hierarchy, which also uses the notation ${\rm co}\mathcal{NP}$; see e.g., [Sto77].
The $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem is deeply connected to the field of proof complexity. Here are some historical notes: In 1979, Cook and Reckhow [CR79] introduced a general definition of a propositional proof system and related it to mainstream complexity theory by pointing out that such a system exists in which all tautologies have polynomial-length proofs if and only if the two complexity classes $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ coincide; see e.g., [CN10]. We refer the reader to the reference [Coo00] for the importance of this research field and to the reference [Kra95] for the motivation of the development of this rich theory. Apart from those, Chapter $10$ of [Pap94] also contains the introductions of the importance of the problem
@@
Likewise, propositional proof complexity is an area of study that has seen a rapid development over the last two decades, which plays as important a role in the theory of feasible proofs as the role played by the complexity of Boolean circuits in the theory of efficient computations; see e.g., the celebrated work [Raz15] by Razborov. In most cases, according to [Raz15], the basic question of propositional proof complexity can be reduced to that given a mathematical statement encoded as a propositional tautology $\psi$ and a class of admissible mathematical proofs formalized as a propositional proof system $P$, what is the minimal possible complexity of a $P$-proof of $\psi$? In other words, propositional proof complexity aims to understand and analyze the computational resources required to prove propositional tautologies, in the same way that circuit complexity studies the resources required to compute Boolean functions. In light of the point of view given in [Raz15], the task of proving lower bounds for strong proof systems like Frege or Extended Frege modulo any hardness assumption in the purely computational world may be almost as interesting; see e.g., [Raz15]. It is worth noting that it has been observed by Razborov [Raz15] that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ implies lower bounds for any propositional proof system. In addition, see e.g., [Raz03] for interesting directions in propositional proof complexity.
@@
Let us return back to the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem for this moment. As a matter of fact, the so-called $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem is the central problem in proof complexity; see e.g., the excellent survey [Kra19]. It formalizes the question of whether or not there are efficient ways to prove the negative cases in our and in many other similar examples. Moreover, the ultimate goal of proof complexity is to show that there is no universal propositional proof system allowing for efficient proofs of all tautologies, which is equivalent to showing that the computational complexity class $\mathcal{NP}$ is not closed under the complement; see e.g., [Kra19].
Particularly, in 1974, Cook and Reckhow showed in [CR74] that there exists a super proof system if and only if $\mathcal{NP}$ is closed under complement; that is, if and only if $\mathcal{NP}={\rm co}\mathcal{NP}$. It is, of course, open whether any superproof system exists. The above result has led to what is sometimes called “Cook’s program” for proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$: prove superpolynomial lower bounds for proof lengths in stronger and stronger proposition proof systems until they are established for all abstract proof systems. Cook’s program, which also can be seen as a program to prove $\mathcal{P}\neq\mathcal{NP}$, is an attractive and plausible approach; unfortunately, it has turned out to be quite hard to establish superpolynomial lower bounds on common proof systems; see e.g., the excellent survey on proof complexity [Bus12].
@@ 2. Preliminaries
Let $\mathbb{N}_{1}=\{1,2,3,\cdots\}$, where $+\infty\notin\mathbb{N}_{1}$, i.e., the set of all positive integers. We also denote by $\mathbb{N}$ the set of natural numbers,333 For the natural number system and $\infty$ (i.e., +$\infty$+ ∞ here), we refer the reader to the interesting textbook [Tao22], in which Tao provides a full and philosophical treatment of them. i.e.,
@@
A ${\rm co}\mathcal{NP}$ machine $M$ is exactly a polynomial-time nondeterministic Turing machine $M$ working under the “there exists no accepting path” condition. We define the language accepted by the ${\rm co}\mathcal{NP}$ machine $M$ to be $\overline{L}(M)$. 444 In short, there is an interesting relationship: ${\rm co}\mathcal{NP}$ machine $M$ accepts the language $\overline{L}(M)$, but polynomial-time nondeterministic Turing machine $M$ rejects $\overline{L}(M)$; polynomial-time nondeterministic Turing machine $M$ accepts the language $L(M)$, but ${\rm co}\mathcal{NP}$ machine $M$ rejects $L(M)$. Namely, for any input $w$, $w\in\overline{L}(M)$ if and only if polynomial-time nondeterministic Turing machine $M$ rejects $w$.
@@ 3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
4. Simulation of All Polynomial-time Nondeterministic Turing Machines
In computational complexity, the most basic approach to show hierarchy theorems uses simulation and diagonalization, because the simulation and diagonalization technique [Tur37, HS65] is a standard method to prove lower bounds on uniform computing models (i.e., the Turing machine model); see for example [AHU74]. These techniques work well in deterministic time and space measures; see e.g., [For00, FS07]. However, they do not work for nondeterministic time, which is not known to be closed under complement at present, i.e., one of the main issues discussed in this paper; hence it is unclear how to define a nondeterministic machine that “does the opposite;” see e.g., [FS07].
For nondeterministic time, we can still do a simulation but can no longer negate the answer directly. In this case, we apply the lazy diagonalization to show the nondeterministic time hierarchy theorem; see e.g., [AB09, SFM78, Zak83]. It is worth noting that Fortnow [For11] developed a much more elegant and simple style of diagonalization to show nondeterministic time hierarchy; see also [FS17]. Generally, lazy diagonalization [Zak83, AB09, SFM78] is a clever application of the standard diagonalization technique [Tur37, HS65, AHU74]. The basic strategy of lazy diagonalization is that for any $n\leq j<m=n^{t(n)}$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $1^{j+1}$, accepting if $M_{i}$ accepts and rejecting if $M_{i}$ rejects. When $j=m$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $x$ deterministically, accepting if $M_{i}$ rejects and rejecting if $M_{i}$ accepts. Since $m=n^{t(n)}$, $M$ has enough time to perform the trivial exponential-time deterministic simulation of $M_{i}$ on input $1^{n}$; what we are doing is deferring the diagonalization step by linking $M$ and $M_{i}$ together on larger and larger inputs until $M$ has an input large enough that it can actually “do the opposite” deterministically; see e.g., [FS07] for details.
@@ 5. Proving That $L_{s}\in\mathcal{NP}$
For any fixed $k\in\mathbb{N}_{1}$, it holds that
@@ 6. Breaking the “Relativization Barrier”
Let us suppose that $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$. In this case, to break the “Relativization Barrier” or to explore the mystery behind the implications of Baker, Gill, and Solovay’s result [BGS75], let us further make the following rational assumptions, which are for comparison with the general point of view about some properties of Turing machines that must be satisfied when applying “diagonalization” techniques (see e.g., p. 73, [AB09]):
@@ A Frege proof system $F$ is an inference system for propositional logic based on
The axiom schemes of a Frege proof system is the following (see e.g., [Bus99]):
@@ 8.1. Lower Bounds for Frege Systems
Initially, propositional proof complexity has been primarily concerned with proving lower bounds (even conditional) for the length of proofs in propositional proof systems, which is extremely interesting and well justified in its own right, with the ultimate goal of settling whether $\mathcal{NP}={\rm co}\mathcal{NP}$ (see e.g., [CR79]). In fact, as mentioned in Section 1, such research directions are called Cook’s Program for separating $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ (see e.g., [Bus12]): Prove superpolynomial lower bounds for proof lengths in stronger and stronger propositional proof systems until they are established for all abstract proof systems. However, our approach is “to do the opposite,” that is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems.
@@ 9. Open Problems
In addition, does the complexity class $\mathcal{BQP}$ (see page 213 of [AB09]) contain the problem Satisfiability? Note that by footnote 1, we have $\mathcal{BQP}\neq\mathcal{NP}$, but this can not exclude that $\mathcal{NP}\subsetneqq\mathcal{BQP}$. If one could show that the problem of Satisfiability is not in $\mathcal{BQP}$, then the relationship between $\mathcal{BQP}$ and $\mathcal{NP}$ would be clear. Currently, we have no additional information about this problem. We tried to think about it but also failed, and we remark that this is also a fascinating open problem in complexity theory.